1/2

Heinrich Heine,  Heinrich Heine,  Heinrich Heine,
The Lorelei  Lorelej’  The Lorelei
tradukita de Mark Twain  tradukita de Leopold Elb  tradukita de L. W. Garnham

An ancient legend of the Rhine Tiu traduko estas poste iom modifita de aliuloj, ne estas la origina traduk-versio de Leopold Elb.

I cannot divine what it meaneth, This haunting nameless pain: A tale of the bygone ages Keeps brooding through my brain:

Ne scias mi, kial mi estas

Malgaja en la kor’. Antikva fabelo restas Por ĉiam en mia memor’.

I do not know what it signifies.

That I am so sorrowful?

A fable of old Times so terrifies, Leaves my heart so thoughtful.

The faint air cools in the gloaming, And peaceful flows the Rhine, The thirsty summits are drinking The sunset’s flooding wine; Vespero jam ekmalheliĝas, Fluadas la Rejn’ en trankvil’, La supro de l’ monto lumiĝas En vespersuna bril’.

The air is cool and it darkens, And calmly flows the Rhine; The summit of the mountain hearkens In evening sunshine line.

The loveliest maiden is sitting High-throned in yon blue air, Her golden jewels are shining, She combs her golden hair; Belega knabino jen tronas Kun ora juvelar’, La belan vizaĝon kronas Mirinde la ora harar’.

The most beautiful Maiden entrances Above wonderfully there, Her beautiful golden attire glances, She combs her golden hair.

She combs with comb that is golden, And sings a weird refrain That steeps in a deadly enchantment The listener’s ravished brain:

Ŝi uzas kombilon el oro Kaj dume kantas ŝi. Mirige tuŝas al koro Sorĉiga la melodi’.

With golden comb so lustrous, And thereby a song sings, It has a tone so wondrous, That powerful melody rings.

…  …  …

2/2

The doomed in his drifting shallop, Is tranced with the sad sweet tone, He sees not the yawing breakers, He sees but the maid alone:

Kaj sopiregon eksentas Ŝipisto dum sia vetur’; Rifaron li ne priatentas, Al monto rigardas li nur.

The shipper in the little ship It effects with woe sad might; He does not see the rocky slip, He only regards dreaded height.

The pitiless billwos engulf him!- So perish sailor and bark; And this, with her baleful singing, Is the Lorelei’s gruesome work.

Mi kredas, ke dronos finfine Ŝipisto kun sia boat’, Kaj tion kaŭzis feine La Loreleja kantad’.

I believe the turbulent waves

Swallow the last shipper and boat; She with her singing craves All to visit her magic moat.

Traduko de la Germana poemo “Die Lorelei” de Heinrich Heine (∗1797-12-13 – †1856-02-17) en la Anglan de Mark Twain.

Arg-2-11 (2003-10-13 04:42:59)

A Tramp Abroad. Vol 1-2. Leibzig: Tauchnitz, 1880 Band I, Mark Twain 1880 (rf. http: // www. loreley. com/ loreley/ marctwai. htm )

Traduko de la Germana poemo “Die Lorelei” de Heinrich Heine (∗1797-12-13 – †1856-02-17) en Esperanton de Leopold Elb (†1912-08.04).

Arg-2-71 (2005-02-03 19:06:30)

La traduko troviĝas sur paĝo 84 de la lernolibro �Post la kurso“ de Wilhelm kaj Hans Wingen, eldonita de Ludwig Pickel en Nurembergo (Nürnberg), Germanio.

Traduko de la Germana poemo “Die Lorelei” de Heinrich Heine (∗1797-12-13 – †1856-02-17) en la Anglan de L. W. Garnham.

Arg-2-14 (2003-10-11 23:04:57)

L.W. Garnham, Bachelor of Arts, LEGENDS OF THE RHINE; mentioned in: A Tramp Abroad. Vol 1-2. Leibzig: Tauchnitz, 1880 Band I, Mark Twain 1880

http://www.poezio.net/version?poem-id=2&version-id=11,71,14